# An ultra-fast intensity interferometry instrument

### v2

@@ Abstract
Context. Intensity interferometry is a reemerging astronomical technique for performing high angular resolution studies at visible wavelengths, benefiting immensely from the recent improvements in (single) photon detection instrumentation. Contrary to direct imaging or amplitude interferometry, intensity interferometry correlates light intensities rather than light amplitudes, circumventing atmospheric seeing limitations at the cost of reduced sensitivity.
Aims. We aim to develop an ultra-fast, single photon counting and highly stable intensity interferometry instrument for 1 m class optical telescopes. The instrument is supposed to record on sky the expected stellar photon rates and reach the temporal coherence times as measured in the laboratory. In addition, all components, especially the photon detection hardware, of the instrument should be easily upgradeable with custom hardware currently being developed.
Methods. The collimated telescope output is spectrally filtered via an ultra narrow band pass of 2 nm at a central wavelength of 405 nm. We use hybrid photon detectors (HPDs) for single photon detection and a constant fraction discriminator (CFD) for signal conditioning. A time to digital converter (TDC) is used for time stamping. The combination of HPDs and CFDs is optimized for large active area and high timing resolution.
Results. We successfully measured photon bunching of three bright A-type stars - Vega, Altair and Deneb at the 1.041.041.04\,1.04m Omicron telescope of C2PU (Centre Pédagogique Planète Univers) at the Calern Observatory in the south of France. In all cases the observed coherence time fits well to both the pre-calculated expectations as well as the values measured in preceding laboratory tests. We obtained the previously estimated photon count rates at the telescope and achieved highly stable coupling of the star light to the detectors.
Conclusions. Utilizing a new class of large area single photon detectors based on multichannel plate amplification, high resolution spatial intensity interferometry experiments are within reach at 111\,1m diameter class telescopes within one night of observation time for bright stars.
Key Words.:
@@ 1 Introduction
The history of stellar intensity interferometry starts with the growing interest in radar technology during the second world war. During that time A.J.F. Siegert looked at the fluctuations in signals finding a relation between the spatio-temporal correlations of the electric fields and of the intensities, today often called the Siegert relation, widely used in intensity interferometry Siegert (1943). Using this relation, Hanbury Brown and Twiss proposed a new type of interferometer to determine the diameter of thermal sources such as stars (Hanbury Brown & Twiss, 1954) by measuring the correlation between photons in two separate beams of light (Hanbury Brown & Twiss, 1956). Shortly afterwards they were able to estimate the angular diameter of Sirius A to $0.0063\,^{\prime\prime}$, a unique precision at that time (Hanbury Brown & Twiss, 2013). Later, the first stellar intensity interferometer at the Narrabri Observatory in New South Wales used the described technique in order to measure the angular diameters of 15 stars (Hanbury Brown et al., 1967a, b).
@@
Recently, stellar intensity interferometry established itself as a complementary technique of current and planed imaging atmospheric Cherenkov telescopes (IACTs) as operation in the presence of bright moonlight is detrimental to observe Cherenkov light but harmless to intensity interferometry measurements (Le Bohec & Holder, 2006). Prominent examples are VERITAS, MAGIC, ASTRI mini array and H.E.S.S. as well as the planned CTA. VERITAS was able to determine the angular diameter of $\beta$ Canis Majoris and $\epsilon$ Orionis with a precision of better than 5% (Abeysekara et al., 2020). The VERITAS Stellar Intensity Interferometer (VSII) was further improved to increase its sensitivity (Kieda et al., 2021a) and since the beginning of 2019 continuing science observations and surveys of stellar targets are carried out(Kieda et al., 2021b). For the IACTs of Major Atmospheric Gamma-Ray Imaging Cherenkov (MAGIC) an optical setup is placed on top of the cameras of the two telescopes (Acciari et al., 2020b). This way MAGIC was able to measure photon correlations (Acciari et al., 2020a) to identify the stellar diameters of three different stars: Adhara ($\epsilon$ CMa), Benetnasch ($\eta$ UMa) and Mirzam ($\beta$ CMa) (Delgado et al., 2021). After hardware modifications MAGIC transitions within seconds to optical interferometry measurements and has determined 22 stellar diameters (Abe et al., 2024). In a similar manner, two telescopes of H.E.S.S were able to record spatio-temporal correlations of three different southern sky stars (Zmija et al., 2024). The ASTRI mini array with its 9 telescopes and baselines between 100 and 700 m is planned to be equipped with fast single photon counting instruments (Zampieri et al., 2022).
@@
Single photon counting detectors have also been used with small optical telescopes to perform intensity interferometry measurements. An example for spatial photon correlation measurements at long baselines is the Asiago Observatory (Italy) with a projected baseline of $\approx 2$ km (Zampieri et al., 2021). With the help of a fixed 1.51.51.5\,1.5m telescope and a 1.01.01.0\,1.0m portable telescope a group from the Université Côte d’Azur was able to observe the extended environment of $\gamma-$Cas at the Calern Observatory in the south of France (Matthews et al., 2023). Elliot Horch et al. (2022) and his group used two out of three portable 0.6 m Dobsonian telescopes equipped with single-photon avalanche diode detectors to observe a correlation peak at the level of $6.76\,\sigma$.
@@ 2.1 Stellar Intensity Interferometry
where for the nominator and denominator averages are taken over time (Labeyrie et al., 2006). $I(t)$ denotes the time dependent intensity and $\gamma^{(1)}(\tau)$ is the first order coherence function also referred to as the complex degree of coherence which can be directly measured in amplitude interferometry. The second part of Eq. 1 is called the Siegert relation that holds for thermal light sources such as stars and connects the second order coherence function to the modulus of the complex degree of coherence (Siegert, 1943). A direct consequence of this relation for thermal light is the effect of ”photon bunching”. Simply speaking, it is more likely to detect a second photon at a detector when a first photon was already recorded within the light coherence time. For example, for a star $\gamma^{(2)}(\tau\rightarrow\infty)=1$ and therefore $\gamma^{(2)}(\tau=0)>\gamma^{(2)}(\tau\rightarrow\infty)$ holds with $\tau$ the delay in time. The width of this intensity correlation excess around zero time delay is proportional to the star’s coherence time $\tau_{c}$. Moreover, the Wiener-Khinchin theorem relates the electric field spectrum $S(\omega)$ of a time-dependent wave to the Fourier transformation of its temporal autocorrelation function (Labeyrie et al., 2006). Hence, we can express the first order coherence function $\gamma^{(1)}(\tau)$ as follows
@@
This convolution leads to a peak with an amplitude $<1$ and a shape similar to the timing jitter. The area under the curve, and thus the coherence time, are however unchanged if $P_{\mathrm{diff}}(t)$ is properly normalized. Note that, when considering only a single polarization mode, the height of the bunching peak is reduced to $\sim\tau_{c}/\Delta t$. Using the relation $\tau_{c}=\lambda^{2}/(c\times\Delta\lambda)$ (Mandel & Wolf, 1995), we can estimate the coherence time of any band pass larger than $\Delta\lambda=0.5\,$nm (FWHM) to $\tau_{c}\ll 1\,$ps. As typical detection systems have a timing jitter around ns to 100100100\,100ps, the detected signal amplitude will be drastically reduced. However, our detection system has a much lower timing jitter of 42.442.442.4\,42.4ps resulting in a better timing resolution, when compared with similar measurements.
@@
where we used the fact that the number of bins $m$ making up the correlation peak is equal to the ratio of the timing resolution $\Delta t$ and the bin size $t_{\mathrm{bin}}$ in the limit $\Delta t\gg\tau_{c}$. For this estimation to hold one has to assume the timing resolution to be twice the measured FWHM resolution to account for full sampling of the bunching peak. It should be pointed out that the count rate $n$ is proportional, whereas the coherence time $\tau_{c}$ is inversely proportional, to the optical bandwidth of the observed light field, hence the S/N is independent of the optical bandwidth to first order. Nevertheless, the usage of a narrow band pass can be beneficial to increase the height of the correlation peak above systematics inherent to the photon detection system. When substituting the conventional definition of the coherence time and spectral flux one arrives at the result of (Hanbury Brown & Twiss, 1957).
@@ 2.3 Experimental setup
The setup used in order to measure temporal correlations is shown in Fig. 1. The light from the telescope enters the setup from the left and is collimated by an achromatic lens with a focus of $-100\,$mm resulting in a beam diameter of 888\,8mm. A short pass dichroic mirror (D1) with a cutoff wavelength of 490490490\,490nm transmits the light used for the intensity correlations. The reflected light is focused via an achromatic lens with a focus of 404040\,40mm onto a ZWO ASI120MM-S camera (ZWO) used for guiding. The transmitted light is passed through a polarizing beam splitter (PBS) which is useful in order to increase the contrast of the bunching peak. Besides it separates light for wavefront analysis without a discernible loss in measurement S/N. The light reflected by the PBS impinges on a self-made Shack-Hartmann sensor. A two lens system is used in order to reduce the beam diameter before a $10\times 10\,$mm microlens array (MLA) with a lens pitch of $150\,\mu$m and a focal length of 5.65.65.6\,5.6mm. A second ZWO is placed in the focal plane of the MLA monitoring the spot diagram, that is the collimation quality of the observed wavefront. This is an important information as an uncollimated beam would degrade the ultra narrow band pass. On the main axis the light is first pre-filtered by a Thorlabs band pass with a central wavelength of 405405405\,405nm and 101010\,10nm bandwidth before passing through a 222\,2nm ultra narrow band pass with a central wavelength of 405405405\,405nm by Alluxa, both shown as rectangles in Fig. 1. This combination is crucial to minimize the possible out-of band transmission. A $50:50$ non-polarizing beam splitter (BS) separates the light onto the two hybrid single photon counting detectors (HPDs) from Becker&Hickl with a quantum efficiency of $22.7\,\%$, see measurement plotted in Appendix C. The signal output from the HPDs is fed through a Constant Fraction Discriminator (CFD) produced by Photonscore to generate NIM pulses. The CFD is then connected to a time to digital converter (TDC) from quTools (quTAG). Due to technical limitations the quTAG introduces a dead time of 404040\,40ns much longer than the 555\,5ns dead time of the detectors. With this rather large dead time start-stop correlations are sufficient for histogramming the photon delay times. In order to account for the 666\,6mm active area of the HPDs, lenses with a focal length of 404040\,40mm slowly converge the beam towards the photocathode, shown as red line inside the HPD. Hence, we are not focusing onto the photocathode but rather only gently decrease the beam diameter to fit the size of the cathode leading to a better beam stability. The CFD threshold is set to a similar value as suggested by the HPD manufacturer, whereas the zero crossing is optimized to obtain a high time resolution of around 303030\,30ps. These two settings lead to dark counts on the order of 505050\,50Hz. The NIM output signal of the CFD then travels to the quTAG through a 111111\,11m and 888\,8m signal cable respectively to perform the correlations. The cable delay of 333\,3m is necessary for the signal to lie inside a nearly shot-noise limited environment, see Appendix B. Post-processing is done via a PC.
@@
A CAD render of the setup and the necessary support structure is shown in Fig. 2. The support structure is mounted directly to the output telescope flange, see Fig. 3. The light from the telescope directly enters the setup having a beam diameter of 888\,8mm when passing through the first lens. The HPDs are movable in the xy-plane, using two stepper motors per detector. Note that due to the high stability of the setup the detector positions once aligned and fixed in the lab, stayed in this position throughout the entire measurement. The setup is designed to fit the constraints of the telescope, these are. weight and dimensions. Counter weights have been used to balance the telescope again after mounting the setup with a weight of approximately 272727\,27kg.
@@
Figure 2: CAD render of the used setup and its support structure. The light enters through the hole on the left. The ring at the front plate is used to center the setup inside the telescope flange, see also Fig. 3. One of the aluminum profiles is transparent to enable a better view onto the setup. The optical components are mounted to the ground plate, see Fig. 1 for a detailed explanation of the optical setup. The hybrid single photon counting detectors (HPDs) are motorized via two stepper motors enabling movement in the xy-plane.
@@
Figure 3: The setup is mounted via $10$ M12 screws to the telescope’s output flange. The light leaves the telescope at the center of its bottom and is directly coupled into our setup hitting the first lens for collimation. The brass cylinders are the necessary counter weights in order to balance the telescope after mounting the approximately 272727\,27kg setup.
@@ 3.1 Lab test
Before testing our setup, we measured the timing resolution of our detection system (TDC, HPDs and CFDs) in the laboratory using a pulsed fs-laser to be 42.442.442.4\,42.4ps (see Appendix  A), which is low compared to typical detection systems that are on the order of 100ps to ns. Together with the filter transmission spectra supplied by the filter manufacturers this measurement allows the calculation of both the expected coherence time and expected shape of the bunching peak using equations Eq 1, 2, 3, and 4. We calculate the expected coherence time to 0.2120.2120.212\,0.212ps, and plot the expected shape of the bunching peak in each of the measurement results shown below. A $2f-2f$ configuration test of our setup, as described in Sec. 2.4, confirmed our expectation and revealed a coherence time of $(0.21\pm 0.02)\,$ps (see Fig. 9).

### v3

# On-sky demonstration of an ultra-fast intensity interferometry instrument utilizing hybrid single photon counting detectors

@@ Abstract
Intensity interferometry is a reemerging astronomical technique for performing high angular resolution studies at visible wavelengths, benefiting immensely from the recent improvements in (single) photon detection instrumentation. Contrary to direct imaging or amplitude interferometry, intensity interferometry correlates light intensities rather than light amplitudes, circumventing atmospheric seeing limitations at the cost of reduced sensitivity. We developed an ultra-fast, single photon counting and highly stable intensity interferometry instrument for 1 m class optical telescopes. The instrument records on sky the expected stellar photon rates and reaches the temporal coherence times as measured in the laboratory. In addition, all components, especially the photon detection hardware, of the instrument are easily upgradeable with custom hardware currently being developed. The collimated telescope output is spectrally filtered via an ultra narrow band pass of 2 nm at a central wavelength of 405 nm. We use hybrid photon detectors (HPDs) for single photon detection and a constant fraction discriminator (CFD) for signal conditioning. A time to digital converter (TDC) is used for time stamping. The combination of HPDs and CFDs is optimized for large active area and high timing resolution. We successfully measured photon bunching of three bright A-type stars - Vega, Altair and Deneb at the 1.041.041.04\,1.04m Omicron telescope of C2PU (Centre Pédagogique Planète Univers) at the Calern Observatory in the south of France. In all cases the observed coherence time fits well to both the pre-calculated expectations as well as the values measured in preceding laboratory tests. We obtained the previously estimated photon count rates at the telescope and achieved highly stable coupling of the star light to the detectors. Utilizing a new class of large area single photon detectors based on multichannel plate amplification, high resolution spatial intensity interferometry experiments are within reach at 111\,1m diameter class telescopes within one night of observation time for bright stars.
keywords:
@@ 1 Introduction
The history of stellar intensity interferometry starts with the growing interest in radar technology during the second world war. During that time A.J.F. Siegert looked at the fluctuations in signals finding a relation between the spatio-temporal correlations of the electric fields and of the intensities, today often called the Siegert relation, widely used in intensity interferometry [1]. Using this relation, Hanbury Brown and Twiss proposed a new type of interferometer to determine the diameter of thermal sources such as stars [2] by measuring the correlation between photons in two separate beams of light [3]. Shortly afterwards they were able to estimate the angular diameter of Sirius A to $0.0063\,^{\prime\prime}$, a unique precision at that time [4]. Later, the first stellar intensity interferometer at the Narrabri Observatory in New South Wales used the described technique in order to measure the angular diameters of 15 stars [5, 6].
@@
Recently, stellar intensity interferometry established itself as a complementary technique of current and planed imaging atmospheric Cherenkov telescopes (IACTs) as operation in the presence of bright moonlight is detrimental to observe Cherenkov light but harmless to intensity interferometry measurements [7]. Prominent examples are VERITAS, MAGIC, ASTRI mini array and H.E.S.S. as well as the planned CTA. VERITAS was able to determine the angular diameter of $\beta$ Canis Majoris and $\epsilon$ Orionis with a precision of better than 5% [8]. The VERITAS Stellar Intensity Interferometer (VSII) was further improved to increase its sensitivity [9] and since the beginning of 2019 continuing science observations and surveys of stellar targets are carried out[10]. For the IACTs of Major Atmospheric Gamma-Ray Imaging Cherenkov (MAGIC) an optical setup is placed on top of the cameras of the two telescopes [11]. This way MAGIC was able to measure photon correlations [12] to identify the stellar diameters of three different stars: Adhara ($\epsilon$ CMa), Benetnasch ($\eta$ UMa) and Mirzam ($\beta$ CMa) [13]. After hardware modifications MAGIC transitions within seconds to optical interferometry measurements and has determined 22 stellar diameters [14]. In a similar manner, two telescopes of H.E.S.S were able to record spatio-temporal correlations of three different southern sky stars [15]. The ASTRI mini array with its 9 telescopes and baselines between 100 and 700 m is planned to be equipped with fast single photon counting instruments [16].
@@
Single photon counting detectors have also been used with small optical telescopes to perform intensity interferometry measurements [17]. An example for spatial photon correlation measurements at long baselines is the Asiago Observatory (Italy) with a projected baseline of $\approx 2$ km [18]. With the help of a fixed 1.51.51.5\,1.5m telescope and a 1.01.01.0\,1.0m portable telescope a group from the Université Côte d’Azur was able to observe the extended environment of $\gamma-$Cas at the Calern Observatory in the south of France [19]. Elliot Horch and his group used two out of three portable 0.6 m Dobsonian telescopes equipped with single-photon avalanche diode detectors to observe a correlation peak at the level of $6.76\,\sigma$ [20].
@@ 2.1 Stellar Intensity Interferometry
where for the nominator and denominator averages are taken over time [21]. $I(t)$ denotes the time dependent intensity and $\gamma^{(1)}(\tau)$ is the first order coherence function also referred to as the complex degree of coherence which can be directly measured in amplitude interferometry. The second part of Eq. 1 is called the Siegert relation that holds for thermal light sources such as stars and connects the second order coherence function to the modulus of the complex degree of coherence [1]. A direct consequence of this relation for thermal light is the effect of ”photon bunching”. Simply speaking, it is more likely to detect a second photon at a detector when a first photon was already recorded within the light coherence time. For example, for a star $\gamma^{(2)}(\tau\rightarrow\infty)=1$ and therefore $\gamma^{(2)}(\tau=0)>\gamma^{(2)}(\tau\rightarrow\infty)$ holds with $\tau$ the delay in time. The width of this intensity correlation excess around zero time delay is proportional to the star’s coherence time $\tau_{c}$. Moreover, the Wiener-Khinchin theorem relates the electric field spectrum $S(\omega)$ of a time-dependent wave to the Fourier transformation of its temporal autocorrelation function [21]. Hence, we can express the first order coherence function $\gamma^{(1)}(\tau)$ as follows
@@
This convolution leads to a peak with an amplitude $<1$ and a shape similar to the timing jitter. The area under the curve, and thus the coherence time, are however unchanged if $P_{\mathrm{diff}}(t)$ is properly normalized. Note that, when considering only a single polarization mode, the height of the bunching peak is reduced to $\sim\tau_{c}/\Delta t$. Using the relation $\tau_{c}=\lambda^{2}/(c\times\Delta\lambda)$ [22], we can estimate the coherence time of any band pass larger than $\Delta\lambda=0.5\,$nm (FWHM) to $\tau_{c}\ll 1\,$ps. As typical detection systems have a timing jitter around ns to 100100100\,100ps, the detected signal amplitude will be drastically reduced. However, our detection system has a much lower timing jitter of 42.442.442.4\,42.4ps resulting in a better timing resolution, when compared with similar measurements.
@@
where we used the fact that the number of bins $m$ making up the correlation peak is equal to the ratio of the timing resolution $\Delta t$ and the bin size $t_{\mathrm{bin}}$ in the limit $\Delta t\gg\tau_{c}$. For this estimation to hold one has to assume the timing resolution to be twice the measured FWHM resolution to account for full sampling of the bunching peak. It should be pointed out that the count rate $n$ is proportional, whereas the coherence time $\tau_{c}$ is inversely proportional, to the optical bandwidth of the observed light field, hence the S/N is independent of the optical bandwidth to first order. Nevertheless, the usage of a narrow band pass can be beneficial to increase the height of the correlation peak above systematics inherent to the photon detection system. When substituting the conventional definition of the coherence time and spectral flux one arrives at the result of [23].
@@ 2.3 Experimental setup
The setup used in order to measure temporal correlations is shown in Fig. 1. The light from the telescope enters the setup from the left and is collimated by an achromatic lens with a focus of $-100\,$mm resulting in a beam diameter of 888\,8mm. This is a major change compared to one of our previous setups used at an optical telescope that featured a deflection mirror at the telescope flange [17]. A short pass dichroic mirror (D1) with a cutoff wavelength of 490490490\,490nm transmits the light used for the intensity correlations. The reflected light is focused via an achromatic lens with a focus of 404040\,40mm onto a ZWO ASI120MM-S camera (ZWO) used for guiding. The transmitted light is passed through a polarizing beam splitter (PBS) which is useful in order to increase the contrast of the bunching peak. Besides it separates light for wavefront analysis without a discernible loss in measurement S/N. The light reflected by the PBS impinges on a self-made Shack-Hartmann sensor. A two lens system is used in order to reduce the beam diameter before a $10\times 10\,$mm microlens array (MLA) with a lens pitch of $150\,\mu$m and a focal length of 5.65.65.6\,5.6mm. A second ZWO is placed in the focal plane of the MLA monitoring the spot diagram, that is the collimation quality of the observed wavefront. This is an important information as an uncollimated beam would degrade the ultra narrow band pass. On the main axis the light is first pre-filtered by a Thorlabs band pass with a central wavelength of 405405405\,405nm and 101010\,10nm bandwidth before passing through a 222\,2nm ultra narrow band pass with a central wavelength of 405405405\,405nm by Alluxa, both shown as rectangles in Fig. 1. This combination is crucial to minimize the possible out-of band transmission. A $50:50$ non-polarizing beam splitter (BS) separates the light onto the two hybrid single photon counting detectors (HPDs) from Becker&Hickl with a quantum efficiency of $22.7\,\%$, see measurement plotted in Appendix C. The signal output from the HPDs is fed through a custom Constant Fraction Discriminator (CFD) produced by Photonscore to generate NIM pulses. This custom CFD enables us to use long cables and a time to digital converter (TDC), the quTAG from quTools, that has a higher timing resolution compared to the time to amplitude converter from our previous work [17]. Due to technical limitations the quTAG introduces a dead time of 404040\,40ns much longer than the 555\,5ns dead time of the detectors. With this rather large dead time start-stop correlations are sufficient for histogramming the photon delay times. In order to account for the 666\,6mm active area of the HPDs, lenses with a focal length of 404040\,40mm slowly converge the beam towards the photocathode, shown as red line inside the HPD. Hence, we are not focusing onto the photocathode but rather only gently decrease the beam diameter to fit the size of the cathode leading to a better beam stability. The CFD threshold is set to a similar value as suggested by the HPD manufacturer, whereas the zero crossing is optimized to obtain a high time resolution of around 303030\,30ps. These two settings lead to dark counts on the order of 505050\,50Hz. The NIM output signal of the CFD then travels to the quTAG through a 111111\,11m and 888\,8m signal cable respectively to perform the correlations. The cable delay of 333\,3m is necessary for the signal to lie inside a nearly shot-noise limited environment, see Appendix B. Post-processing is done via a PC.
@@
A CAD render of the setup and the necessary support structure is shown in Fig. 2(a). The support structure is mounted directly to the output telescope flange, see Fig. 2(b). The light from the telescope directly enters the setup having a beam diameter of 888\,8mm when passing through the first lens. The HPDs are movable in the xy-plane, using two stepper motors per detector. Note that due to the high stability of the setup the detector positions once aligned and fixed in the lab, stayed in this position throughout the entire measurement. The setup is designed to fit the constraints of the telescope, these are. weight and dimensions. Counter weights have been used to balance the telescope again after mounting the setup with a weight of approximately 272727\,27kg.
@@
Figure 2: (a) CAD render of the used setup and its support structure. The light enters through the hole on the left. The ring at the front plate is used to center the setup inside the telescope flange, see also Fig. 2(b). One of the aluminum profiles is transparent to enable a better view onto the setup. The optical components are mounted to the ground plate, see Fig. 1 for a detailed explanation of the optical setup. The hybrid single photon counting detectors (HPDs) are motorized via two stepper motors enabling movement in the xy-plane. (b) The setup is mounted via $10$ M12 screws to the telescope’s output flange. The light leaves the telescope at the center of its bottom and is directly coupled into our setup hitting the first lens for collimation. The brass cylinders are the necessary counter weights in order to balance the telescope after mounting the approximately 272727\,27kg setup.
@@ 3.1 Lab test
Before testing our setup, we measured the timing resolution of our detection system (TDC, HPDs and CFDs) in the laboratory using a pulsed fs-laser to be 42.442.442.4\,42.4ps (see Appendix  A), which is low compared to typical detection systems that are on the order of 100ps to ns. Together with the filter transmission spectra supplied by the filter manufacturers this measurement allows the calculation of both the expected coherence time and expected shape of the bunching peak using equations Eq 1, 2, 3, and 4. We calculate the expected coherence time to 0.2120.2120.212\,0.212ps, and plot the expected shape of the bunching peak in each of the measurement results shown below. A $2f-2f$ configuration test of our setup, as described in Sec. 2.4, confirmed our expectation and revealed a coherence time of $(0.21\pm 0.02)\,$ps (see Fig. 6(b)).